| 12/19/2004 | ||||
| 0:22 | ESC asked us to check pervious uploading (PMF010) | HP | ||
| So, we send "SCI_SW_LOGFILE_SEND" | ||||
| >>recevied 20041219-051852.sys | ||||
| Using grep, confirmed pmf010.00x 8 all files are uploaded and assembled | ||||
| 1:29 | Next command window will start at 2:21(EST) | YS | ||
| ES and Sonny call to send command to undo recording on the flash disk | ||||
| I asked ESC to send commands using Iridium. | ||||
| They said that they are going to set up soon. | ||||
| Buffer is going up tp 5% and started making pds files. | ||||
| 1:36 | ESC said they are connected through Iridium and ready to send commands. So, I sent following commands | YS | ||
| SCI_SNIO_CONTROL 12, 0x08 | ||||
| SCI_SNIO_CONTROL 13, 0x44 | ||||
| Also, send commands for reslove making pds files. | ||||
| SCI_SW_OPMODE 0 | YS | |||
| SCI_SNIO_CONTROL 21,0 | ||||
| Execute script daq_restart.stol | ||||
| 1:57 | start making pds files again. | YS | ||
| 2:25 | Commanding window opened | YS | ||
| SCI_SW_OPMODE 0 | ||||
| SCI_SNIO_CONTROL 21,0 | ||||
| Execute script daq_restart.stol | ||||
| ESC said they are going to continue uploading fragment. | ||||
| 4:09 | SCI_SW_OPMODE 0 | HP | ||
| SCI_SNIO_CONTROL 21,0 | ||||
| Execute script daq_restart.stol | ||||
| 4:12 | SCI_SW_OPMODE 0 | HP | ||
| SCI_SNIO_CONTROL 21,0 | ||||
| Execute script daq_restart.stol | ||||
| 4:28 | According to Sonny's email, send a command | YS | ||
| SCI_SW_CUSTOM | ||||
| And compared a file received and another file attached at Sonny's email. Results are different so, I stopped here. | ||||
| 4:30 | ESC said we don't have TDRSS data flow window for 30 minutes. | YS | ||
| No more data flow and housekeeping updates. | ||||
| 5:42 | SCI_SW_OPMODE 0 | HP | ||
| SCI_SNIO_CONTROL 21,0 | ||||
| Execute script daq_restart.stol | ||||
| 5:53 | SCI_SW_OPMODE 0 | YS | ||
| SCI_SNIO_CONTROL 21,0 | ||||
| Execute script daq_restart.stol | ||||
| 5:55 | Housekeeping does not update and data does not come.. | YS | ||
| ESC said they are connected to us but they don't see any data flowing. Since there was no housekeeping update, I think, it looks like connection problem. So, I asked ESC to disconnect data flow and reconnect it. ESC said they are getting their data from CDM. | ||||
| 6:06 | Now we are in the between commanding windows until 6:46 | YS | ||
| 6:46 | Still data and houdekeeping does not come | |||
| ESC said they are receving their CDM information | ||||
| 7:10 | Talked to Opher about this and he suggest to ask ESC to send ping from CDM to SFC. | |||
| ESC said that they make CDM send a packet to SFC but there was no response. | ||||
| 7:30 | Talked Larry, he suggest to send command to reset software. | |||
| I asked ESC to send commands using Iridium. | ||||
| 7:40 | SCI_SW_DAQ_RESET | |||
| SCI_SW_RESTART | ||||
| 11:00 | Still data and houdekeeping does not come | JH | ||
| We
decided to turn off HDS MB02 of the power to watch about the change of current with CDM monitoring. |
||||
| 11:30 | sent command "SCI_SW_NOOP", but failed | JH | ||
| Ping is blocked | ||||
| 11:42 | SCI_SW_NOOP -> This command went through to ESC and CDM | JH | ||
| 11:43 | SCI_HDS_DC_POWER 0x02, OFF | |||
| HDS MB02(before: 1.85~1.87, after: 1.78~1.86) | ||||
| 11:46 | SCI_HDS_DC_POWER 0x02, ON | |||
| It's changed just a little bit, so we tried again about CAL_MB05 | ||||
| 12:02 | SCI_CAL_DC_POWER 0x05, OFF | JH | ||
| CAL_MB05(before: 2.07, after:2.05) | ||||
| maybe this command didn't turn it off. | ||||
| 20:21 | asked ESC to shut down flight computer. | |||
| 21:19 | We tried to turn off the flight computer. However because it is so dangerous method, we waited for EunSuk`s permittion, | |||
| 23:00 | ES approved that sfc will be rebooted if we have confidence to proceed to turning on processes. | |||
| 23:30 | sent command "SCI_SW_NOOP" | |||